OpenStudy (raffle_snaffle):

If a sum is worth $550 today and it will be worth $768 two years from now, how much was it worth one year ago?

OpenStudy (raffle_snaffle):

@mathmale

OpenStudy (raffle_snaffle):

Don't just give me the answer. Help me walk to the solution.

OpenStudy (anonymous):

\[\frac{ 550 + 768 }{ 2 } = 659\] Gives you the value after 1 year. I'm assuming this is reflexive so then... 659 - initial value = 109$ increase or decrease per year. Take the initial value of 550 - 109. This is the price from the year before the current year.

OpenStudy (raffle_snaffle):

How does (550+768)/2 give you the value after 1 year

OpenStudy (raffle_snaffle):

are you taking the average or does 2 represent 2 years?

OpenStudy (r.hamza17):

He is using 2 for the 2 years..

OpenStudy (raffle_snaffle):

Okay I think I get it with a little more analyzing.

OpenStudy (r.hamza17):

Where do you need more analyzing?

OpenStudy (raffle_snaffle):

So over 1 year the amount increased by 109. So in 2 years the increase went up twice 2*109 = 218

OpenStudy (raffle_snaffle):

I got it. Lol Thanks for your help, both of you.